# Introducing NewCluster: the first half of the history of a high-resolution cluster simulation

### v1

@@ 3.4 Gas duty cycles and AGN feedback
Gas motion is intrinsically complex, governed not only by secular hydrodynamics but also by energetic astrophysical events such as SN explosions and AGN outflows. NewCluster is performed via AMR technique, an Eulerian approach that excels at capturing shocks and discontinuities during the feedback processes (Teyssier, 2002; O’Shea et al., 2005; Hubber et al., 2013). However, despite this strength, AMR does not track the motion of individual gas cells, limiting its ability to follow Lagrangian gas trajectories, which is important for understanding the impact of feedback processes (Bellovary et al., 2013; Choi et al., 2024). In galaxy evolution studies, this makes it difficult to distinguish between inflowing and outflowing gas, which is crucial for understanding the cold gas supply and the impact of feedback on star formation.
To overcome this limitation, we apply Monte-Carlo tracer particles that follow the hydrodynamics of gas cells (Genel et al., 2013; Cadiou et al., 2019). Cadiou et al. (2019) demonstrated that Monte-Carlo tracers exhibit better performance than velocity-advected models in tracing the gas motion and spatial distribution in cosmological simulations. The simulation includes 267,997,944 tracer particles in total. These tracer particles probabilistically move between cells and stars, and massive black hole particles based on their gas mass exchange events, such as flux between cells, star formation, mass ejection from feedback, and accretion. Tracer particles provide valuable opportunities by allowing the tracking of actual gas motions via particle IDs within the simulation framework. We found that the spatial distribution of gas tracers closely resembles the density distribution of gas cells within the simulation. However, we note that the discrete number of tracer particles cannot perfectly reproduce the true transport of gas mass, a limitation shared with the full Lagrangian approach.
In Fig. 8, we show the trajectories of gas tracer particles within a galaxy as an example of their use in tracking gas dynamics. The central galaxy ($M_{*}=2.2\times 10^{12}\,\,M_{\sun}$ in $M_{\rm host}=4.63\times 10^{13}\,\,M_{\sun}$) experiences a violent AGN outburst at $z=1.6$. We first selected sample gas tracers that were initially located within $R_{\rm 90}$ of the central galaxy and further classified them into three categories based on their trajectories before and after $500\,\mathrm{Myr}$:
@@
    Inflow: The time when the tracer is farthest from the central MBH occurs before reaching its pericentre, and its maximum distance exceeds $10\,R_{\rm 90}$.
@@
    Outflow: The tracer reaches a maximum distance greater than $6\,R_{\rm 90}$ and continues to move away from the centre for at least $400\,{\rm Myr}$.
@@
    Recycle: The tracer travels outward beyond $2.5\,R_{\rm 90}$ but later returns within $1.5\,R_{\rm 90}$.
We show the smoothed trajectories of individual gas tracer particles, classified into each category and colour-coded accordingly. The background displays the gas temperature. Since the target galaxy is moving from left to right, most of the inflowing gas (blue) originates from the right side of the figure. However, it is clearly visible that some gas inflow also originates from the bottom, exhibiting a narrowly-channeled feature that indicates gas brought in by an infalling satellite galaxy. We note that the theoretical study reported that the most stream influx comes from one or few dominant streams (Danovich et al., 2012). Outflowing tracers (red) rapidly escape from the galactic centre along the bipolar AGN jet axis, which is aligned with the galaxy’s rotation axis. They also show highly turbulent motion after reaching the low-density circumgalactic region. Recycled tracers (green) do not display a strongly preferential direction, in contrast to inflow or outflow tracers. The presence of recycled gas demonstrates that a significant amount of gas can escape from the galactic disc and later return, consistent with the galactic fountain scenario (Bregman, 1980; Melioli et al., 2008).
Tracer particles are directly applicable to investigate how AGN feedback ejects galactic gas and regulates central star formation. Moreover, as our simulation is set in a cluster environment, it is also suited to studying how gas in satellite galaxies is stripped in the intracluster medium (ICM) by ram pressure or how their gas supply is halted. Tracer particles also provide a means to study the MBH duty cycle (Martini & Weinberg, 2001; Choi et al., 2024). Since we track the state transition of tracer particles to gas, star, and MBH tracers, we can distinguish the gas consumed by accretion from that expelled by feedback.
@@ 3.5 Ram-pressure stripping of gas
Cluster environments host the most significant gas removal processes observed in galaxies. The interaction between the hot ICM and the ISM of galaxies moving at velocities of $\sim 1000\,{\rm km\,s^{-1}}$ exerts strong pressure, estimated as $P_{\rm ram}/k_{\rm B}\sim 10^{5-6}\,{\rm K}\,{\rm cm}^{-3}$ (Jung et al., 2018; Yun et al., 2019; Boselli et al., 2021), following the formulation of Gunn & Gott (1972). This external pressure often exceeds the internal gravitational restoring force of galaxies, leading to the stripping of their gas components. Observations of gas trails including jellyfish phenomenon (e.g., Kenney et al., 2004; Chung et al., 2009; Jaffé et al., 2018) provide evidence for this mechanism. The resulting gas loss effectively quenches star formation (SF), often on short timescales although a temporal enhancement of SF activity occurs due to compression of ISM (e.g., Crowl & Kenney, 2006; Vulcani et al., 2018; Roberts & Parker, 2020). Consequently, strong ram pressure, in combination with other environmental mechanisms, is now recognised as a driver of the high fraction of passive galaxies observed in clusters (e.g., Contini et al., 2020; Rhee et al., 2020; Oman et al., 2021).
@@
An important next step is to understand how complex the interactions between the ICM and ISM truly are, and how gas is stripped from galaxies in a manner that reproduces the observed morphologies. One effective approach is the use of idealised wind-tunnel simulations, where a uniform ICM wind is applied to a model galaxy to mimic ram pressure stripping (e.g., Bekki, 2009; Tonnesen & Bryan, 2012; Lee et al., 2020; Choi et al., 2022). These simulations offer very high spatial and temporal resolution, enabling detailed studies of gas dynamics under controlled conditions. However, a key limitation is that the ram pressure strength is often assumed to be constant, which does not reflect the time-varying conditions experienced by galaxies in live halos (c.f, Tonnesen, 2019). Recent studies suggest that such variations in ram pressure may play a role in shaping both the gas content and morphology of galaxies (e.g., Samuel et al., 2023; Zhu et al., 2024). Therefore, it is essential to incorporate realistic ram pressure profiles when studying its effects on galaxies.
@@
Cosmological hydrodynamic simulations provide a complementary framework for studying the impact of ram pressure on galaxy evolution. For example, Jung et al. (2018) used the YZiCS simulation (Choi & Yi, 2017) to investigate gas removal in clusters, finding that strong ram pressure near cluster centres is a dominant driver of gas deficiency in satellite galaxies. Several other studies employing cosmological simulations (e.g., Lotz et al., 2019; Yun et al., 2019; Rohr et al., 2023) have similarly concluded that ram pressure plays a central role in driving environmental quenching in clusters. However, these cosmological simulations often lack the spatial resolution necessary to capture the complexity of ICM-ISM interactions, particularly the multiphase or clumpy nature of ISM, which may be more resilient to stripping. As a result, there is significant potential for overestimating or overlooking ram pressure effects in cosmological simulations, especially when small-scale physical processes fall below the resolution limit.
@@
The NewCluster simulation bridges the gap between idealised and cosmological approaches to studying ram pressure stripping. Its spatial resolution is sufficient to capture high-density gas components—possibly corresponding to the molecular phase of ISM—that are expected to be less sensitive to ram pressure (e.g., Tonnesen & Bryan, 2009). In addition, the mass resolution enables the modeling of ram pressure effects on dwarf galaxies with $M_{*}>10^{7-9}\,M_{\odot}$, which is a mass range that remains largely inaccessible to most large-scale simulations. Although NewCluster does not extend to the local Universe, it captures the high-redshift regime ($z>0.8$), a period when ram pressure was already a significant environmental process (e.g., Simons et al., 2020). As such, the simulation offers an opportunity to investigate the early formation of gas-poor galaxies in dense environments and to shed light on the onset of environmental quenching in nascent clusters.
@@
The upper panel of Fig. 9 illustrates the gas distributions within the main cluster in NewCluster at $z=1$, where the gas column density is colour-coded and the virial radius of the main cluster is shown as the grey circle. Although the cluster has not yet reached the canonical mass range for local clusters ($M_{\rm vir}>10^{14}\,M_{\odot}$), numerous galaxies display extended gas tails, indicating that ram pressure effects are already active at $z=1$. To evaluate whether ram pressure is sufficiently strong to influence galaxy evolution, we compute its value for individual galaxies surrounding clusters. The volume-weighted mean ram pressure is measured using ICM gas cells, following the methodology of the ICM gas selection in Rhee et al. (2024):
@@
where $\rho_{i}$, ${v}_{i}$, and $V_{i}$ are the density, velocity with respect to the galaxy, and the volume of $i$-th gas ICM cell, respectively. The bottom left panel of Fig. 9 presents the radial profile of ram pressure across the galaxy sample. A clear decline in ram pressure with increasing distance is seen, primarily reflecting the decreasing ICM density at larger radii. The scatter in this trend is likely driven by variations in galaxy velocities. In addition, galaxy–galaxy interactions can also induce strong ram pressure. For example, galaxies experiencing strong ram pressure at around $D_{\rm 3D}=250$ or $450\,{\rm kpc}$ are satellites undergoing significant ram pressure from the gas disc of a massive galaxy. Notably, galaxies located near the cluster core experience ram pressure values on the order of $P_{\rm ram}/k_{\rm B}\sim 10^{6}\,{\rm K}\,{\rm cm}^{-3}$, consistent with estimates for local clusters (Jung et al., 2018; Yun et al., 2019; Boselli et al., 2021).
@@ 3.6 Dust properties
Figure 10: An example galaxy extinction curve and corresponding mock observation. The extinction curve is shown in the first column of the figure. The total extinction curve (grey) is presented along with the contributions from graphite (cyan) and silicate (magenta) grains, each shown with two different size distributions (dotted lines for small grains and dashed lines for large grains). The MW (Fitzpatrick & Massa, 2007) extinction, SMC (Pei, 1992) extinction, and Calzetti attenuation curves are also included for comparison. The grain size distributions for different grain types are shown in the sub-panel, with the MRN (Mathis et al., 1977) slope indicated by a yellow dashed line. The spatial distribution of the DTM is shown in the second column. The corresponding mock image, created using SDSS g, r, and i bands for RGB, is shown in the last column.
@@
Numerical simulations have recently begun to follow dust evolution in a cosmological context on the fly, including dust formation, destruction, and the evolution of chemical composition and grain size (Davé et al., 2019; Jones et al., 2024; Choban et al., 2025). This approach goes well beyond the simpler post-processing methods that map dust onto the metal mass content of the gas. Since NewCluster tracks the evolution of carbonaceous and silicate grains separately, it enables a detailed examination of the distinct behaviours of different dust species as they respond to dynamic environmental changes such as mergers, AGN feedback, and starbursts on short timescales. We also trace the evolution of grain size. Under the two-size approximation, small and large grains form and evolve through different but interrelated processes. By combining grain size and chemical composition information, we can investigate the evolution of the intrinsic dust extinction curve and generate realistic mock observations, even at high redshifts.
@@
Figure 10 shows the dust properties of a sample spiral galaxy at $z=0.8$. The left panel shows the intrinsic extinction curve based on the average composition of dust inside $3R_{\rm 50}$ using Mie theory and grain optical constants from Weingartner & Draine (2001). The total extinction curve is shown in grey, and the cyan and magenta lines represent the carbon-based and silicon-based grain species, respectively. The contributions from small and large grains are shown with different line styles (dotted and dashed lines for small and large grains, respectively), and the grain size distribution is also presented in a sub-panel in the upper right corner.
@@
The Milky Way (MW) average extinction curve (Fitzpatrick & Massa, 2007), the Small Magellanic Cloud (SMC) extinction curve (Pei, 1992), and the Calzetti attenuation curve for starburst galaxies (Calzetti et al., 2000) are also given for reference. We note that, unlike the MW and SMC extinction curves, the Calzetti curve represents an attenuation law that encapsulates both absorption and scattering effects in a mixed star–dust geometry (for reference, see Narayanan et al. 2018, which illustrates how Calzetti-like attenuation laws can emerge from different extinction curves depending on the geometry between stars and dust). The sample NewCluster galaxy shows a UV-to-optical slope similar to the Calzetti attenuation curve, but notably with a clear 2175Å bump as shown in the MW curve produced by small carbonaceous grains.
@@
The spatial distribution of the dust-to-metal mass ratio (DTM) is also presented in the middle panel. The mean DTM inside $3R_{\rm 50}$ is $\sim 0.4$. We find that the DTM of the dust lanes following the spiral arms (see the RGB composite mock image in the right panel, generated using SKIRT with the SDSS $u$-, $g$-, and $r$-bands mapped to blue, green, and red, respectively) is nearly 0.5 or above, but shows various values over a wide range. We notice a large cavity of DTM ratio in the centre and small ones in the spiral arms. They are the sites of enhanced star formation, where SN shocks destroy a large fraction of the dust. The central DTM cavity is of particular significance, because it seriously affects the apparent light distribution of galaxies and thus morphology determination at high redshifts (Byun et al. 2025, submitted).

### v2

@@ 3.4 Gas duty cycles and AGN feedback
Gas motion is intrinsically complex, governed not only by secular hydrodynamics but also by energetic astrophysical events such as SN explosions and AGN outflows. NewCluster is performed via AMR technique, an Eulerian approach that excels at capturing shocks and discontinuities during the feedback processes (2002A&A...385..337T; O'Shea2005; Hubber2013). However, despite this strength, AMR does not track the motion of individual gas cells, limiting its ability to follow Lagrangian gas trajectories, which is important for understanding the impact of feedback processes (Bellovary2013; Choi2024). In galaxy evolution studies, this makes it difficult to distinguish between inflowing and outflowing gas, which is crucial for understanding the cold gas supply and the impact of feedback on star formation.
To overcome this limitation, we apply Monte-Carlo tracer particles that follow the hydrodynamics of gas cells (Genel2013; Cadiou2019). Cadiou2019 demonstrated that Monte-Carlo tracers exhibit better performance than velocity-advected models in tracing the gas motion and spatial distribution in cosmological simulations. The simulation includes 267,997,944 tracer particles in total. These tracer particles probabilistically move between cells and stars, and massive black hole particles based on their gas mass exchange events, such as flux between cells, star formation, mass ejection from feedback, and accretion. Tracer particles provide valuable opportunities by allowing the tracking of actual gas motions via particle IDs within the simulation framework. We found that the spatial distribution of gas tracers closely resembles the density distribution of gas cells within the simulation. However, we note that the discrete number of tracer particles cannot perfectly reproduce the true transport of gas mass, a limitation shared with the full Lagrangian approach.
In Fig. 8, we show the trajectories of gas tracer particles within a galaxy as an example of their use in tracking gas dynamics. The central galaxy ($M_{*}=2.2\times 10^{12}\,\,M_{\sun}$ in $M_{\rm host}=4.63\times 10^{13}\,\,M_{\sun}$) experiences a violent AGN outburst at $z=1.6$. We first selected sample gas tracers that were initially located within $R_{\rm 90,3D}$ of the central galaxy and further classified them into three categories based on their trajectories before and after $500\,\mathrm{Myr}$:
@@
    Inflow: The time when the tracer is farthest from the central MBH occurs before reaching its pericentre, and its maximum distance exceeds $10\,R_{\rm 90,3D}$.
@@
    Outflow: The tracer reaches a maximum distance greater than $6\,R_{\rm 90,3D}$ and continues to move away from the centre for at least $400\,{\rm Myr}$.
@@
    Recycle: The tracer travels outward beyond $2.5\,R_{\rm 90,3D}$ but later returns within $1.5\,R_{\rm 90,3D}$.
We show the smoothed trajectories of individual gas tracer particles, classified into each category and colour-coded accordingly. The background displays the gas temperature. Since the target galaxy is moving from left to right, most of the inflowing gas (blue) originates from the right side of the figure. However, it is clearly visible that some gas inflow also originates from the bottom, exhibiting a narrowly-channeled feature that indicates gas brought in by an infalling satellite galaxy. We note that the theoretical study reported that the most stream influx comes from one or few dominant streams (Danovich2012). Outflowing tracers (red) rapidly escape from the galactic centre along the bipolar AGN jet axis, which is aligned with the galaxy’s rotation axis. They also show highly turbulent motion after reaching the low-density circumgalactic region. Recycled tracers (green) do not display a strongly preferential direction, in contrast to inflow or outflow tracers. The presence of recycled gas demonstrates that a significant amount of gas can escape from the galactic disc and later return, consistent with the galactic fountain scenario (Bregman1980; Melioli2008).
Tracer particles are directly applicable to investigate how AGN feedback ejects galactic gas and regulates central star formation. Moreover, as our simulation is set in a cluster environment, it is also suited to studying how gas in satellite galaxies is stripped in ICM by ram pressure or how their gas supply is halted. Tracer particles also provide a means to study the MBH duty cycle (Martini2001; Choi2024). Since we track the state transition of tracer particles to gas, star, and MBH tracers, we can distinguish the gas consumed by accretion from that expelled by feedback.
@@ 3.5 Ram-pressure stripping of gas
Cluster environments host the most significant gas removal processes observed in galaxies. The interaction between the hot ICM and the ISM of galaxies moving at velocities of $\sim 1000\,{\rm km\,s^{-1}}$ exerts strong pressure, estimated as $P_{\rm ram}/k_{\rm B}\sim 10^{5-6}\,{\rm K}\,{\rm cm}^{-3}$ (Jung18; Yun19; Boselli21), following the formulation of GG72. This external pressure often exceeds the internal gravitational restoring force of galaxies, leading to the stripping of their gas components. Observations of gas trails including jellyfish phenomenon (e.g., Kenney04; Chung09; Jaffe18) provide evidence for this mechanism. The resulting gas loss effectively quenches star formation (SF), often on short timescales although a temporal enhancement of SF activity occurs due to compression of ISM (e.g., Crowl06; Vulcani18; Roberts20). Consequently, strong ram pressure, in combination with other environmental mechanisms, is now recognised as a driver of the high fraction of passive galaxies observed in clusters (e.g., Contini20; Rhee20; Oman21).
@@
An important next step is to understand how complex the interactions between the ICM and ISM truly are, and how gas is stripped from galaxies in a manner that reproduces the observed morphologies. One effective approach is the use of idealised wind-tunnel simulations, where a uniform ICM wind is applied to a model galaxy to mimic ram pressure stripping (e.g., Bekki09; Tonnesen12; Lee20; Choi22). These simulations offer very high spatial and temporal resolution, enabling detailed studies of gas dynamics under controlled conditions. However, a key limitation is that the ram pressure strength is often assumed to be constant, which does not reflect the time-varying conditions experienced by galaxies in live halos (c.f, Tonnesen19). Recent studies suggest that such variations in ram pressure may play a role in shaping both the gas content and morphology of galaxies (e.g., Samuel23; Zhu24). Therefore, it is essential to incorporate realistic ram pressure profiles when studying its effects on galaxies.
@@
Cosmological hydrodynamic simulations provide a complementary framework for studying the impact of ram pressure on galaxy evolution. For example, Jung18 used the YZiCS simulation (CY17) to investigate gas removal in clusters, finding that strong ram pressure near cluster centres is a dominant driver of gas deficiency in satellite galaxies. Several other studies employing cosmological simulations (e.g., Lotz19; Yun19; Rohr23) have similarly concluded that ram pressure plays a central role in driving environmental quenching in clusters. However, these cosmological simulations often lack the spatial resolution necessary to capture the complexity of ICM-ISM interactions, particularly the multiphase or clumpy nature of ISM, which may be more resilient to stripping. As a result, there is significant potential for overestimating or overlooking ram pressure effects in cosmological simulations, especially when small-scale physical processes fall below the resolution limit.
@@
The NewCluster simulation bridges the gap between idealised and cosmological approaches to studying ram pressure stripping. Its spatial resolution is sufficient to capture high-density gas components—possibly corresponding to the molecular phase of ISM—that are expected to be less sensitive to ram pressure (e.g., Tonnesen09). In addition, the mass resolution enables the modeling of ram pressure effects on dwarf galaxies with $M_{*}>10^{7-9}\,M_{\odot}$, which is a mass range that remains largely inaccessible to most large-scale simulations. Although NewCluster does not extend to the local Universe, it captures the high-redshift regime ($z>0.8$), a period when ram pressure was already a significant environmental process (e.g., Simons20). As such, the simulation offers an opportunity to investigate the early formation of gas-poor galaxies in dense environments and to shed light on the onset of environmental quenching in nascent clusters.
@@
The upper panel of Fig. 9 illustrates the gas distributions within the main cluster in NewCluster at $z=1$, where the gas column density is colour-coded and the virial radius of the main cluster is shown as the grey circle. Although the cluster has not yet reached the canonical mass range for local clusters ($M_{\rm vir}>10^{14}\,M_{\odot}$), numerous galaxies display extended gas tails, indicating that ram pressure effects are already active at $z=1$. To evaluate whether ram pressure is sufficiently strong to influence galaxy evolution, we compute its value for individual galaxies surrounding clusters. The volume-weighted mean ram pressure is measured using ICM gas cells, following the methodology of the ICM gas selection in Rhee24:
@@
where $\rho_{i}$, $\@vec{v}_{i}$, and $V_{i}$ are the density, velocity with respect to the galaxy, and the volume of $i$-th gas ICM cell, respectively. The bottom left panel of Fig. 9 presents the radial profile of ram pressure across the galaxy sample. A clear decline in ram pressure with increasing distance is seen, primarily reflecting the decreasing ICM density at larger radii. The scatter in this trend is likely driven by variations in galaxy velocities. In addition, galaxy–galaxy interactions can also induce strong ram pressure. For example, galaxies experiencing strong ram pressure at around $D_{\rm 3D}=250$ or $450\,{\rm kpc}$ are satellites undergoing significant ram pressure from the gas disc of a massive galaxy. Notably, galaxies located near the cluster core experience ram pressure values on the order of $P_{\rm ram}/k_{\rm B}\sim 10^{6}\,{\rm K}\,{\rm cm}^{-3}$, consistent with estimates for local clusters (Jung18; Yun19; Boselli21).
@@ 3.6 Dust properties
Figure 10: An example galaxy extinction curve and corresponding mock observation. The extinction curve is shown in the first column of the figure. The total extinction curve (grey) is presented along with the contributions from graphite (cyan) and silicate (magenta) grains, each shown with two different size distributions (dotted lines for small grains and dashed lines for large grains). The MW (Fitzpatrick&Massa2007) extinction, SMC (Pei1992) extinction, and Calzetti attenuation curves are also included for comparison. The grain size distributions for different grain types are shown in the sub-panel, with the MRN (Mathis1977) slope indicated by a yellow dashed line. The spatial distribution of the DTM is shown in the second column. The corresponding mock image, created using SDSS g, r, and i bands for RGB, is shown in the last column.
@@
Numerical simulations have recently begun to follow dust evolution in a cosmological context on the fly, including dust formation, destruction, and the evolution of chemical composition and grain size (Dave2019; Jones2024; Choban2025). This approach goes well beyond the simpler post-processing methods that map dust onto the metal mass content of the gas. Since NewCluster tracks the evolution of carbonaceous and silicate grains separately, it enables a detailed examination of the distinct behaviours of different dust species as they respond to dynamic environmental changes such as mergers, AGN feedback, and starbursts on short timescales. We also trace the evolution of grain size. Under the two-size approximation, small and large grains form and evolve through different but interrelated processes. By combining grain size and chemical composition information, we can investigate the evolution of the intrinsic dust extinction curve and generate realistic mock observations, even at high redshifts.
@@
Figure 10 shows the dust properties of a sample spiral galaxy at $z=0.8$. The left panel shows the intrinsic extinction curve based on the average composition of dust inside $3R_{\rm 50,3D}$ using Mie theory and grain optical constants from Weingartner2001. The total extinction curve is shown in grey, and the cyan and magenta lines represent the carbon-based and silicon-based grain species, respectively. The contributions from small and large grains are shown with different line styles (dotted and dashed lines for small and large grains, respectively), and the grain size distribution is also presented in a sub-panel in the upper right corner.
@@
The Milky Way (MW) average extinction curve (Fitzpatrick&Massa2007), the Small Magellanic Cloud (SMC) extinction curve (Pei1992), and the Calzetti attenuation curve for starburst galaxies (Calzetti2000) are also given for reference. We note that, unlike the MW and SMC extinction curves, the Calzetti curve represents an attenuation law that encapsulates both absorption and scattering effects in a mixed star–dust geometry (for reference, see Narayanan2018, which illustrates how Calzetti-like attenuation laws can emerge from different extinction curves depending on the geometry between stars and dust). The sample NewCluster galaxy shows a UV-to-optical slope similar to the Calzetti attenuation curve, but notably with a clear 2175Å bump as shown in the MW curve produced by small carbonaceous grains.
@@
The spatial distribution of the dust-to-metal mass ratio (DTM) is also presented in the middle panel. The mean DTM inside $3R_{\rm 50,3D}$ is $\sim 0.4$. We find that the DTM of the dust lanes following the spiral arms (see the RGB composite mock image in the right panel, generated using SKIRT with the SDSS $u$-, $g$-, and $r$-bands mapped to blue, green, and red, respectively) is nearly 0.5 or above, but shows various values over a wide range. We notice a large cavity of DTM ratio in the centre and small ones in the spiral arms. They are the sites of enhanced star formation, where SN shocks destroy a large fraction of the dust. The central DTM cavity is of particular significance, because it seriously affects the apparent light distribution of galaxies and thus morphology determination at high redshifts (2025ApJ...992...92B).